OpenStudy (anonymous):

1 + cos(12x) = _____. A. 2sin2(6x) B. 2sin2(12x) C. 2cos2(12x) D. 2cos2(6x)

OpenStudy (anonymous):

I believe it is 1.814180971. I literally just typed this equation in my graphing calculator and it came up with this.

OpenStudy (mathstudent55):

Use this identity: \( \cos 2\theta = 2\cos^2 \theta - 1 \)